Find a general solution for the differential equation

(D^2 - 3D +2) = (e^2x)/(1+e^2x)
1
Expert's answer
2020-05-27T13:03:49-0400

the auxiliary equation f the given differential equation is:

D2-3D+2=0

D=2,1

y(x)= c1e2x+c2ex


particular solution by method of variation of parameters:


let u(x)=e2x

v=ex

f= "\frac{e^{2x}}{1+e^{2x}}"


then, wronskian W(u,v)="\begin{vmatrix}
 e^{2x} & e^{x} \\
 2e^{2x} & e^{x}
\end{vmatrix}" = -e3x


then, the particular solution is:

"=-u\int{\frac{v.fdx}{W}}+v\int{\frac{u.fdx}{W}}"


"= -e^{2x}\int{\frac{e^x . e^{2x}dx}{-e^{3x}(1+e^{2x})}}+ e^{x}\int{\frac{e^{2x} . e^{2x}dx}{-e^{3x}(1+e^{2x})}}"


"=e^{2x} \int{\frac{1}{1+e^{2x}}}dx - e^x\int{\frac{e^x}{1+e^{2x}}}dx"


"= e^{2x}.ln|sin(tan^{-1}.e^x)|-e^x tan^{-1}e^x"


so the solution is:

"y=c_1e^{2x}+c_2e^x+e^{2x}.ln|sin(tan^{-1}.e^x)|-e^x tan^{-1}e^x"